# Learning from Videos for 3D World: Enhancing MLLMs with 3D Vision Geometry Priors

### v2

@@ Abstract
Previous research has investigated the application of Multimodal Large Language Models (MLLMs) in understanding 3D scenes by interpreting them as videos. These approaches generally depend on comprehensive 3D data inputs, such as point clouds or reconstructed Bird’s-Eye View (BEV) maps. In our research, we advance this field by enhancing the capability of MLLMs to understand and reason in 3D spaces directly from video data, without the need for additional 3D input. We propose a novel and efficient method, the Video-3D Geometry Large Language Model (VG LLM). Our approach employs a 3D visual geometry encoder that extracts 3D prior information from video sequences. This information is integrated with visual tokens and fed into the MLLM. Extensive experiments have shown that our method has achieved substantial improvements in various tasks related to 3D scene understanding and spatial reasoning, all directly learned from video sources. Impressively, our 4B model, which does not rely on explicit 3D data inputs, achieves competitive results compared to existing state-of-the-art methods, and even surpasses the Gemini-1.5-Pro in the VSI-Bench evaluations.
@@ 1 Introduction
Extensive experiments have been conducted on various 3D scene understanding and spatial reasoning tasks, where the model accepts video input. These 3D scene understanding tasks include 3D visual grounding [7], 3D dense captioning [8], and 3D video object detection [46]. For spatial reasoning tasks, we evaluate our model on VSI-Bench [51], CV-Bench [41], and BLINK [19]. The experiments show that our fine-tuned 4B model outperforms larger spatial-enhanced models by a substantial margin. The results uncover several interesting findings: (1) Without explicit dense 3D inputs, our approach outperforms many leading 3D input-based models, underscoring its effective 3D geometric understanding. (2) By implicitly modeling inter-frame correspondences within the visual representation, our model has learned strong egocentric-allocentric transformation capabilities, leading to significant improvements of precision by 9.1 and F1 by 8.2 on 3D video object detection. (3) On tasks that require complex spatial reasoning skills, i.e., VSI-Bench [51], our 4B model attains an impressive average score of 46.1%, surpassing even the best proprietary model, Gemini-1.5-Pro [40]. This highlights the great utility of 3D geometry modeling in broader scenarios.
@@ 3.2.1 Applying VG LLM to 3D Scene Understanding
We follow the recent work [61, 66] that decomposes the 3D dense caption task into two phases, i.e., detecting 3D object proposals with an off-the-shelf detector and generating object descriptions based on object coordinates. Specifically, given the 3D scene $\{I_{1},I_{2},\cdots,I_{n}\}$, we prompt our model to “describe the object located at $(x,y,z)$ in detail”, where $(x,y,z)$ is the box center of the object to be described. This setup ensures a fair comparison with prior work and effectively evaluates the model’s ability to understand positional and spatial relationships.
@@ 4 Experiments
In this section, we first present the implementation details of our model, followed by an evaluation of our model’s performance on 3D scene understanding tasks in Section 4.1, including 3D visual grounding, 3D dense captioning, and 3D multi-view detection. Next, we provide a comprehensive comparison with state-of-the-art methods on spatial reasoning benchmarks and generic multimodal benchmarks in Section 4.2. 3D scene understanding emphasizes 3D perception, while spatial reasoning focuses on interpreting and reasoning about spatial relationships in video. We train two models for 3D scene understanding and spatial reasoning tasks separately for fair comparison. Finally, we conduct ablation studies to reveal the effectiveness of integrating 3D geometry into MLLMs in Section 4.3.
Implementation Details. Our model is built upon Qwen2.5-VL-3B [2], integrated with VGGT-1B [43] as the 3D geometry encoder. We trained our model for one epoch on a mixed dataset, detailed as followed in this section, and employed Adam with a batch size of 16 and a warmup ratio of 0.03. During the warmup phase, the learning rate was gradually increased to 1e-5 before linearly decaying to 0. In each training step, a batch was randomly sampled from a single source from the mixed dataset. Please refer to the appendix for more details.
@@ 4.1.2 Results on 3D Visual Grounding
We follow SPAR [57] to perform the refinement process by matching the predicted bounding box with the pre-detected object proposals.
Model 3D Scene Input Acc@0.25 Acc@0.5
@@
SPAR [57] 48.8 (31.9) 43.1 (12.4)
VG LLM (Ours) 53.5 (36.4) 47.5 (11.8)
Table 1: The quantitative results on ScanRefer. The content in “()” indicates results without proposal refinement.
Model 3D Scene Input C@0.5$\uparrow$ B-4@0.5$\uparrow$ M@0.5$\uparrow$ R@0.5$\uparrow$
@@
VG LLM (Ours) 78.6 40.9 28.6 62.4
Table 2: The performance on Scan2Cap. VG LLM does not use the ground truth 3D scene information.
Figure 2: Visualization of 3D visual grounding results. The ground truth and prediction are masked in blue and green, respectively. The predicted boxes are directly generated by our model without the refinement process.
VG LLM shows excellent visual grounding performance without 3D information.
We show the performance in ScanRefer in Table 2. Our model obtains an Accuracy of 53.5 at an IoU of 0.25, outperforming SPAR’s accuracy of 48.8 by 4.7%. Even without the proposal refinement process, our model also delivers an Accuracy of 36.4, which surpasses SPAR’s 31.9 by 4.5%. This demonstrates that 3D visual grounding can be effectively approached in a video grounding manner.
Visualization for 3D visual grounding.
Figure 2 presents a visualization of the 3D visual grounding results. To visualize the predicted and ground truth bounding boxes, we transform their bounding boxes to the respective image coordinate system. Our method effectively handles spatial relationships between objects (e.g., away, opposite, next to), accurately identifying the corresponding frame index and generating the 3D-oriented bounding box.
@@ 4.1.3 Results on 3D Dense Captioning
Pure-vision solution works surprisingly well for 3D dense captioning.
We present the performance on Scan2Cap in Table 2. Despite not utilizing 3D camera parameters or explicit depth information, our model achieves competitive results, surpassing the previous state-of-the-art LEO by 6.2 in CIDEr. This indicates that the 3D geometric information implicitly encoded within the visual features extracted by VGGT is sufficiently informative for indoor scene understanding, indicating the potential of pure-vision solutions for 3D dense captioning.
@@ 4.1.4 Results on 3D Video Object Detection
Model Frames chair cabinet table bin couch bed bathtub toilet 20 Common Classes
@@
Qwen2.5-VL-3B 4 37.7 10.2 35.0 23.1 39.0 64.8 32.4 68.8 32.6 27.9 30.0
+ Visual Geometry (VG LLM) 49.7 13.1 41.3 39.2 44.6 71.2 33.5 83.4 41.7 35.7 38.2
@@
Qwen2.5-VL-3B 6 32.8 7.8 31.3 20.9 32.2 58.8 36.5 66.1 27.8 24.1 25.7
+ Visual Geometry (VG LLM) 41.6 12.4 39.8 33.1 45.0 70.2 33.8 80.6 39.7 34.0 36.4
@@
Figure 3: Visualization of 3D video object detection results.
As shown in Table 3, we follow the monocular detection of EmbodiedScan [46] to evaluate the performance for categories that are common in daily life. For evaluation, we report the per-class F1 score as well as the average precision, recall, and F1 score.
VG LLM significantly outperforms the baselines.
Qwen2.5-VL [2], when fine-tuned on detection data, achieves decent performance on common object categories. Furthermore, incorporating visual geometry [43] yields a substantial improvement across all evaluation metrics. Notably, the average F1 in the 4-frame setting exhibits a dramatic increase of 8.2, rising from 30.0 to 38.2. This improvement can be attributed to the model’s enhanced ability to understand the egocentric-allocentric transformation between frames through the utilization of visual geometry. Consequently, the model can retrieve more objects that were not visible in the first frame. Figure 2 shows the visualization on 3D visual grounding.
Integrating 3D geometry enhances the model’s robustness to the number of frames.
Table 3 shows that our method, although trained on 4-frame sequences, maintains strong performance when evaluated on longer sequences during inference. In contrast, the baseline’s performance drops noticeably, highlighting the greater robustness of our approach to variations in the number of frames.
@@ 4.2 Spatial Reasoning Benchmarks
Model Avg. Numerical Answer Multiple-Choice Answer
@@ 4.2.1 Setting
To fully leverage the 3D knowledge inherent in the 3D vision geometry encoder, we train our model on a dataset sampled from SPAR-7M [57] and the LLaVA-Hound split of the LLaVA-Video-178K [60]. In our work, we sample only 234K and 63K data points from SPAR-7M and the LLaVA-Hound split of LLaVA-Video-178K, respectively, which constitute 3% and 25% of the original datasets.
We first evaluate our method on a video spatial reasoning benchmark, VSI-Bench [51], which evaluates the egocentric-allocentric transformation and relational reasoning capabilities of MLLMs. Then we test our model on two image spatial-related benchmarks i.e., CV-Bench [41] and BLINK (spatial subset) [19]. CV-Bench [41] assesses 2D understanding through spatial relationships and object counting, while its 3D evaluation focuses on depth ordering and relative distance perception. The spatial subset of BLINK [19] includes three tasks requiring 3D spatial perception abilities, namely relative depth, spatial reasoning, and multi-view reasoning.
@@ 4.2.2 Results on Spatial Reasoning Tasks
Model Depth (%) Spatial (%) Multi-View (%) Avg. (%)
Proprietary Models (API)
GPT-4V [35] 60.0 72.7 55.6 62.8
GPT-4o [24] 74.2 69.2 59.4 67.6
Claude-3-OPUS [1] 47.6 58.0 56.4 54.0
Open-source Models
CogVLM [47] 52.4 63.6 54.1 56.7
LLaVA-v1.5-13B-xtuner [33] 54.0 69.9 44.4 56.1
Yi-VL-34B [53] 50.0 71.3 44.4 55.2
LLaVA-v1.6-34B [33] 67.7 74.8 62.4 68.3
SAT-LLaVA-Video-7B [38] 66.1 73.4 48.1 62.6
VG LLM-4B (Ours) 79.8 71.3 54.1 68.4
Table 6: The comparison with the SOTA methods on the validation split of BLINK (spatial subsets).
Our method achieves state-of-the-art performance on VSI-Bench.
We present the detailed comparison with the top-tier models in Table 4. As shown in the table, our model achieves state-of-the-art performance with an average score of 46.1 among all compared models, even surpassing the best-performing proprietary model, Gemini-1.5 Pro. Remarkably, our model secures the highest score in counting and room size estimation, reaching 66.4% and 56.3%, respectively. This highlights the model’s strong capability in understanding and interpreting scene configurations.
Our model generalizes well across distinct data sources.
To investigate the generalization ability of our method, we also evaluate our approach on the spatial reasoning benchmarks with different data sources. CV-Bench is constructed by repurposing the traditional CV datasets (i.e., ADE20K, COCO, and Omni3D) to a vision-centric MLLM benchmark. As illustrated in Table 6, our method achieves the highest accuracy on 3D tasks at 91.3%. BLINK (spatial subset) comprises a wide range of image sources, including Wild6D [54], VSR [32], and Depth in the Wild [13]. Moreover, we present the results on BLINK in Table 6. Our 4B model achieves the highest score on relative depth at 79.8, surpassing GPT-4o by 5.6. These results demonstrate our model can also generalize well across the spatial reasoning benchmarks with out-of-the-domain data sources.
A small amount of training data suffices for strong performance.
Unlike SPAR [57], which relies on a massive 2 million samples for pre-training, our approach demonstrates that leveraging just a small amount of data (3% of SPAR-7M) can achieve strong performance by effectively capturing the embedded visual geometry. This finding highlights the importance of explicit modeling 3D visual geometry in MLLMs. Furthermore, reducing the amount of fine-tuning data helps mitigate distribution drift from the original MLLM, thereby enhancing its generalization capability.
@@ 4.2.3 Results on Generic Multimodal Benchmarks
[…]
Qwen2.5-VL-3B 60.1 60.9 46.6 62.2 77.3
VG LLM-4B 57.3 59.8 50.4 63.4 74.2
Table 7: Comparison of model performance on generic multimodal benchmarks.
@@
4.3 Ablation Study
Spatial Signal ScanRefer Scan2Cap 3D Video Detection
@@ 4.3 Ablation Study
None 49.9 (31.9) 43.8 (9.3) 58.0 36.3 32.6 27.9 30.0
Pred Camera Info. 50.0 (32.1) 43.9 (9.9) 56.8 36.3 33.3 28.1 30.3
Pred Depth Info. 49.7 (32.3) 43.7 (9.7) 57.1 36.1 32.1 27.3 29.3
Pred (Depth + Camera) Info. 49.8 (32.6) 43.7 (9.8) 58.2 36.6 31.7 27.1 29.1
Pred Point Info. 49.7 (31.7) 43.8 (9.6) 67.7 38.2 33.0 28.5 30.4
3D Visual Geometry 53.5 (36.4) 47.5 (11.8) 78.6 40.9 41.7 35.7 38.2
Table 8: Ablation study of the effects of 3D visual geometry modeling. All models are fine-tuned using the same training data.
To demonstrate that the improvement in spatial reasoning and 3D scene understanding in MLLMs is genuinely driven by visual geometry modeling, we conducted an ablation study, as shown in Table 8. Besides the visual geometry feature adopted in our model, VGGT can also directly predict the camera poses (Pred Camera Info.), the depth maps (Pred Depth Info.), and the point maps (Pred Point Info.) of each image frame. For these predicted spatial signals, we adopt a two-layer MLP to transform them and add to the corresponding 2D visual token. All models are trained and evaluated following the datasets and benchmarks in Sec. 4.1.
@@
VSI-Bench CV-Bench
Abs. Dist. Room Size Obj. Size Route Plan Depth Distance
Qwen2.5-VL-3B 19.7 25.4 17.2 29.4 74.5 66.0
    + finetune 31.9 47.8 54.3 27.8 85.3 87.8
    + Visual Geometry (VG LLM) 36.6 56.3 55.2 30.4 93.7 89.8
Table 9: Ablation study on spatial reasoning tasks.
Visual geometry improves spatial reasoning, specifically in measurement estimation.
We also ablate the effect of finetuning and visual geometry in Table 9. The figure illustrates that fine-tuning on our mixed dataset leads to a substantial performance gain in measurement estimation (e.g., absolute distance, room size). After incorporating the visual geometry into the model architecture, it further improves absolute distance and room size by 4.7% and 8.5%, respectively.

### v3

@@ Abstract
Previous research has investigated the application of Multimodal Large Language Models (MLLMs) in understanding 3D scenes by interpreting them as videos. These approaches generally depend on comprehensive 3D data inputs, such as point clouds or reconstructed Bird’s-Eye View (BEV) maps. In our research, we advance this field by enhancing the capability of MLLMs to understand and reason in 3D spaces directly from video data, without the need for additional 3D input. We propose a novel and efficient method called the Video-3D Geometry Large Language Model (VG LLM). Our approach utilizes a 3D visual geometry encoder to extract 3D prior information from video sequences. This information is then integrated with visual tokens and input into the MLLM. Extensive experiments have shown that our method has achieved substantial improvements in various tasks related to 3D scene understanding and spatial reasoning, all directly learned from video sources. Impressively, our 4B model, which does not rely on explicit 3D data inputs, achieves competitive results compared to existing state-of-the-art methods, and even surpasses the Gemini-1.5-Pro in the VSI-Bench evaluations.
@@ 1 Introduction
Extensive experiments have been conducted on various 3D scene understanding and spatial reasoning tasks, where the model accepts video input. These 3D scene understanding tasks include 3D visual grounding [7], 3D dense captioning [8], and 3D video object detection [45]. For spatial reasoning tasks, we evaluate our model on VSI-Bench [50] and CV-Bench [40]. The experiments show that our fine-tuned model outperforms larger spatial-enhanced models by a substantial margin. The results uncover several interesting findings: (1) Without explicit dense 3D inputs, our approach outperforms many leading 3D input-based models, underscoring its effective 3D geometric understanding. (2) By implicitly modeling inter-frame correspondences within the visual representation, our 8B model has learned strong egocentric-allocentric transformation capabilities, leading to significant improvements of precision by 11.9% and F1 by 10.7% on 3D video object detection. (3) On tasks that require complex spatial reasoning skills, i.e., VSI-Bench [50], our 4B model attains an impressive average score of 47.3%, surpassing even the best proprietary model, Gemini-1.5-Pro [39]. Furthermore, our 8B model sets a new state-of-the-art performance with a score of 50.7%. This highlights the great utility of 3D geometry modeling in broader scenarios.
@@ 3.2.1 Applying VG LLM to 3D Scene Understanding
We follow the recent work [58, 63] that decomposes the 3D dense captioning task into two phases, i.e., detecting 3D object proposals with an off-the-shelf detector and generating object descriptions based on object coordinates. Specifically, given a frame sequence $\{I_{1},I_{2},\cdots,I_{n}\}$, we prompt our model to “describe the object located at $(x,y,z)$ in detail”, where $(x,y,z)$ is the box center of the object to be described. This setup ensures a fair comparison with prior work and effectively evaluates the model’s ability to understand positional and spatial relationships.
@@ 4 Experiments
In this section, we first present the implementation details of our model, followed by an evaluation of our model’s performance on 3D scene understanding tasks in Section 4.1, including 3D visual grounding, 3D dense captioning, and 3D multi-view detection. Next, we provide a comprehensive comparison with state-of-the-art methods on spatial reasoning benchmarks and generic multimodal benchmarks in Section 4.2. 3D scene understanding emphasizes 3D perception, while spatial reasoning focuses on interpreting and reasoning about spatial relationships in video. We train two models for 3D scene understanding and spatial reasoning tasks separately for fair comparison. Finally, we conduct a comprehensive analysis in Section 4.3 to reveal the effectivess of our model’s core aspects: the integration of 3D geometry, the feature fusion strategy, and the data composition.
Implementation and Training Details. Our models are built upon two sizes of Qwen2.5-VL—3B and 7B [2], and integrated with VGGT-1B [42] as the 3D geometry encoder. We trained our models for one epoch on a mixed dataset, detailed as followed in this section, and employed Adam with a batch size of 64 and a warmup ratio of 0.03. During the warmup phase, the learning rate was gradually increased to 1e-5 before linearly decaying to 0. In each training step, a batch was randomly sampled from a single source from the mixed dataset. During training, the MLLM’s visual encoder, the integrated 3D geometry encoder, and the multimodal connector are frozen, while the MLLM backbone remains unfrozen. All experiments were conducted on 8 H100 80G GPUs. For the 4B and 8B models, the training took 9 and 12 hours for 3D scene understanding, and 7 and 9 hours for spatial reasoning, respectively.
@@ 4.1.1 Setting
Evaluation Metrics.
For ScanRefer [7], we report the accuracy at IoU thresholds of 0.25 and 0.5. For Scan2Cap [8], we use CIDEr (C), BLEU-4 (B-4), METEOR (M), and ROUGE (R) scores. The “@0.5” suffix indicates that these metrics are computed only for objects detected with an IoU of 0.5 or higher against the ground truth. Lastly, for 3D video object detection, we report the accuracy, recall, and F1 score for 20 common object classes at an IoU threshold of 0.25.
@@ 4.1.2 Results on 3D Visual Grounding
Model
3D Scene
Input
Acc@0.25 Acc@0.5
@@
SPAR [54] 31.9 (48.8) 12.4 (43.1)
VG LLM-4B (Ours) 36.4 (53.5) 11.8 (47.5)
VG LLM-8B (Ours) 41.6 (57.6) 14.9 (50.9)
Table 1: The quantitative results on ScanRefer. The content in “()” indicates results with proposal refinement [^2] For proposal refinement, we compare the predicted box against all proposals detected by Mask3D and select the one with the highest IoU..
Model
3D Scene
Input
[…]
@@
VG LLM-4B (Ours) 78.6 40.9 28.6 62.4
VG LLM-8B (Ours) 80.0 41.5 28.9 62.6
Table 2: The performance on Scan2Cap. VG LLM captions 3D object proposals using RGB cues only.
Figure 2: Qualitative results for 3D visual grounding. The ground truth and prediction are masked in blue and green, respectively. The predicted boxes are directly generated by our model without the refinement process.
In addition to the directly predicted bounding boxes, we also follow SPAR [54] to calculate the refined results with a proposal refinement process33footnotemark: 3.
VG LLM demonstrates strong 3D grounding capabilities from monocular RGB videos.
The performance on ScanRefer is shown in Table 2. Our 8B model achieves an accuracy of 41.6% at an IoU threshold of 0.25, significantly outperforming SPAR’s 31.9% by 9.7 percentage points. Moreover, with the integration of a proposal refinement technique, Acc@0.25 increases substantially to 57.6%, making our results highly competitive with state-of-the-art methods like Video-3D LLM. This demonstrates that 3D visual grounding can be effectively approached in a video grounding manner.
Qualitative results for 3D visual grounding.
Figure 2 illustrates the qualitative results on the 3D visual grounding task. For visualization purposes, the predicted and ground-truth 3D bounding boxes are projected onto their respective 2D image planes. Our method effectively handles spatial relationships between objects (e.g., away, opposite, next to), accurately identifying the corresponding frame index and generating the 3D-oriented bounding box.
@@ 4.1.3 Results on 3D Dense Captioning
Our model processes 3D object proposals generated by an off-the-shelf 3D object detector.
Once these objects are localized, the model generates descriptions using only the RGB images and the center coordinates of the detected objects, without requiring explicit depth or geometric data as additional inputs. This approach achieves a score of 80.0 C@0.5 and 41.5 B-4@0.5, comparable to previous state-of-the-art methods such as LLaVA-3D and Video-3D LLM. These results validate the effectiveness of our model for 3D dense captioning when conditioned on 3D proposals.
@@ 4.1.4 Results on 3D Video Object Detection
Model chair cabinet table bin couch bed bathtub toilet 20 Common Classes
@@
4-Frame Setting
Qwen2.5-VL-3B 37.7 10.2 35.0 23.1 39.0 64.8 32.4 68.8 32.6 27.9 30.0
+ Visual Geometry (VG LLM-4B) 49.7 13.1 41.3 39.2 44.6 71.2 33.5 83.4 41.7 35.7 38.2
@@
Qwen2.5-VL-7B 41.2 11.6 36.5 30.2 41.1 68.2 36.6 68.7 34.6 31.0 32.5
+ Visual Geometry (VG LLM-8B) 54.0 17.1 46.5 39.8 47.0 74.1 42.1 82.5 43.4 39.6 41.2
$\triangle$ Improvement +12.8 +5.5 +10.0 +9.7 +5.9 +5.9 +5.5 +13.8 +8.8 +8.6 +8.7
6-Frame Setting
Qwen2.5-VL-3B 32.8 7.8 31.3 20.9 32.2 58.8 36.5 66.1 27.8 24.1 25.7
+ Visual Geometry (VG LLM-4B) 41.6 12.4 39.8 33.1 45.0 70.2 33.8 80.6 39.7 34.0 36.4
@@
Figure 3: Qualitative results for 3D video object detection.
As shown in Table 3, we follow the monocular detection of EmbodiedScan [45] to evaluate the performance for categories that are common in daily life.
3D geometry significantly boosts cross-frame detection performance.
Qwen2.5-VL [2], upon being fine-tuned on detection data, achieves promising performance on common object categories. Incorporating 3D geometry [42] yields a substantial improvement across all evaluation metrics. In the 4-frame setting, this approach elevates the average F1 score for the 4B model by 8.2% (from 30.0% to 38.2%) and for the 8B model by 8.7% (from 32.5% to 41.2%). This performance gain is attributed to the model’s improved ability to comprehend the egocentric-allocentric transformation between frames through the utilization of 3D geometry. Figure 3 provides qualitative results on the 3D video object detection task.
Integrating 3D geometry enhances the model’s robustness to variations in frame count.
Table 3 shows that our method, although trained on 4-frame sequences, maintains strong performance when evaluated on longer sequences during inference. In contrast, the baseline’s performance drops noticeably, underscoring the robustness of our approach to variations in the number of frames.
@@ 4.2 Spatial Reasoning Benchmarks
Model
Avg. Numerical Answer Multiple-Choice Answer
@@ 4.2.1 Setting
To fully leverage the 3D knowledge inherent in the 3D visual geometry encoder, we train our model on a dataset sampled from SPAR-7M [54] and the LLaVA-Hound split of the LLaVA-Video-178K [57]. In our work, we sample only 234K and 63K data points from SPAR-7M and the LLaVA-Hound split of LLaVA-Video-178K, respectively, which constitute 3% and 25% of the original datasets.
We first evaluate our method on a video spatial reasoning benchmark, VSI-Bench [50], which evaluates the egocentric-allocentric transformation and relational reasoning capabilities of MLLMs. Then we test our model on an image spatial-related benchmark i.e., CV-Bench [40]. CV-Bench [40] assesses 2D understanding through spatial relationships and object counting, while its 3D evaluation focuses on depth ordering and relative distance perception.
@@
Evaluation Metrics.
For VSI-Bench, we adopt accuracy for multiple-choice tasks and Mean Relative Accuracy (MCA) for numerical tasks. MCA calculates the average accuracy across a range of confidence thresholds, where a prediction is considered correct if its relative error is within a specified threshold. For CV-Bench and other generic multimodal benchmarks, we report accuracy.
@@ 4.2.2 Results on Spatial Reasoning Tasks
Benchmark
Qwen2.5-VL-3B
VG LLM-4B
Qwen2.5-VL-7B
VG LLM-8B
[…]
BLINK [19] 46.6 48.9 55.9 51.5
[…]
Table 6: Comparison of model performance on generic multimodal benchmarks. To ensure a fair comparison, we evaluate Qwen2.5-VL with the resolution and sampling rate matched to those of VG LLM.
VG LLM achieves state-of-the-art performance on VSI-Bench.
A detailed comparison with leading models is presented in Table 4. As shown in the table, our 4B model achieves an impressive average accuracy of 47.3%, outperforming all competitors, including the leading proprietary model, Gemini-1.5 Pro. Futhermore, our 8B model sets a new state-of-the-art performance with an average accuracy of 50.7%. These results underscore the model’s strong capability in understanding and reasoning from monocular videos.
Our model demonstrates generalization across distinct data sources.
To access the generalization capability of our method, we evaluate our approach on a spatial reasoning benchmark with different data sources, CV-Bench. This benchmark is constructed by repurposing the traditional CV datasets (i.e., ADE20K, COCO, and Omni3D) to a vision-centric MLLM benchmark. As illustrated in Table 6, our method achieves the highest accuracy on 3D tasks at 91.1%. These results demonstrate our model can also generalize well across the spatial reasoning benchmarks with out-of-the-domain data sources.
@@ 4.2.3 Results on Generic Multimodal Benchmarks
4.3 Analysis
4.3.1 3D Scene Understanding
Type
ScanRefer Scan2Cap 3D Video Detection
@@ 4.3 Analysis
No Additional Info. (Baseline) 31.9 (49.9) 9.3 (43.8) 58.0 36.3 32.6 27.9 30.0
Cross-Attn (1 Layer) 33.7 (51.3) 10.7 (45.2) 74.7 40.1 38.0 33.6 35.4
Cross-Attn (3 Layers) 34.4 (51.3) 10.5 (45.2) 75.7 40.2 38.5 44.0 35.4
Concat+MLP 27.7 (47.0) 6.8 (41.3) 75.7 39.9 37.1 32.5 34.4
Add (Ours) 36.4 (53.5) 11.8 (47.5) 78.6 40.9 41.7 35.7 38.2
Pred Camera Info. 32.1 (50.0) 9.9 (43.9) 56.8 36.3 33.3 28.1 30.3
Pred Depth Info. 32.3 (49.7) 9.7 (43.7) 57.1 36.1 32.1 27.3 29.3
Pred Point Info. 31.7 (49.7) 9.6 (43.8) 67.7 38.2 33.0 28.5 30.4
Pred (Depth + Camera) Info. 32.6 (49.8) 9.8 (43.7) 58.2 36.6 31.7 27.1 29.1
Latent 3D Geometry (Ours) 36.4 (53.5) 11.8 (47.5) 78.6 40.9 41.7 35.7 38.2
Table 7: Ablation study of the effects of 3D visual geometry modeling. All models are fine-tuned using the same training data and built upon Qwen2.5-VL-3B.
In this section, we conduct a further analysis on the feature fusion strategies and the types of added signals. To investigate the effect of feature fusion strategies, we have experimented with several feature fusion strategies, including 1) “Cross-Attn”, which consists of multiple blocks of cross-attention modules and MLPs with skip connections. 2D visual tokens serve as queries, while 3D visual tokens serve as keys and values. 2D positional embeddings are added to both the queries and keys; 2) “Concat+MLP” first concatenates the 2D and 3D visual tokens along the feature dimension, followed by an MLP to transform them into the text embedding space; 3) “Add” is the strategy employed in our paper, which directly adds the 2D and 3D visual tokens at a patch level.
Besides the visual geometry feature adopted in our model, VGGT can also directly predict the camera poses (Pred Camera Info.), the depth maps (Pred Depth Info.), and the point maps (Pred Point Info.) of each image frame. For these predicted spatial signals, we adopt a two-layer MLP to transform them and add to the corresponding 2D visual token. All models are trained and evaluated following the datasets and benchmarks in Section 4.1.
Directly adding the geometry features obtains the best results among all compared feature fusion strategies.
As Table 7 illustrates, the “Cross-Attn” strategy significantly improves performance over the baseline. Moreover, increasing the number of layers to three yields even greater performance gains. While “Concat+MLP” performs worse than the baseline on ScanRefer, it outperforms it on Scan2Cap and 3D video object detection. Nevertheless, “Add” surpasses all other comparison methods despite its simplicity.
@@
4.3.2 Spatial Reasoning
# Data Model VSI-Bench CV-Bench (3D)
Obj. Count Abs. Dist. Obj. Size Room Size Rel. Dist. Rel. Dir. Route Plan Appr. Order Avg Depth Distance Avg
1 $\emptyset$ Qwen2.5-VL-7B 25.4 10.4 36.4 29.1 38.2 37.9 30.9 27.5 29.5 84.5 76.5 80.5
2 S1 Qwen2.5-VL-7B 68.6 36.0 57.9 59.7 45.8 38.9 30.4 60.2 49.8 87.0 87.2 87.1
3 VG LLM-8B 67.9 37.7 58.6 62.0 46.6 40.7 32.4 59.2 50.7 92.3 89.8 91.1
4 S2 Qwen2.5-VL-7B 71.3 48.3 68.5 65.5 65.1 77.5 40.2 16.7 56.6 66.8 74.5 70.7
5 VG LLM-8B 71.7 53.8 68.8 62.1 63.8 83.0 44.3 18.4 58.2 66.8 73.3 70.1
6 S1+S2 Qwen2.5-VL-7B 70.2 51.0 69.8 64.4 67.0 79.1 45.4 31.6 59.8 87.0 83.8 85.4
7 VG LLM-8B 71.4 56.8 69.0 69.1 67.9 83.2 47.4 32.5 62.2 92.0 89.5 90.8
Table 8: Ablation study on spatial reasoning tasks. $\emptyset$ indicates testing in a zero-shot setting. S1 is the training data detailed in Section 4.2.1, and S2 refers to the VLM-3R dataset.
We then investigate the effects of fine-tuning and visual geometry in Table 8. The table illustrates that fine-tuning on our mixed dataset (S1) leads to a substantial performance gain in measurement estimation (rows 1 vs. 2). In addition to the data used in this paper (S1), the VLM-3R dataset [16] (S2), which is specially curated for the task types in VSI-Bench, is also incorporated in our analysis. After incorporating the 3D geometry into the model architecture, the results show consistent gains with the different data compositions, i.e., S1, S2 and S1+S2.